ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 06A15; 06A23 Zapatrin, Roman, E & F University, St. Petersburg, Russia Duality and set representations for partially ordered sets A representation theory for arbitrary partially ordered sets based on a Stonelike duality is built. It is shown how Stone and MacLaren representations are incorporated in the suggested framework. Some results on completions are presented.